OpenStudy (anonymous):

the number of real solutions of x-1/(x^2-25) = 5 - 1/(x^2-25) is?

OpenStudy (ash2326):

@Yahoo! did you try this?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

i think i must cross multiply

OpenStudy (ash2326):

Yeah, do that;D

OpenStudy (anonymous):

naah, easier thing,

OpenStudy (anonymous):

so it will be cubic equation!

OpenStudy (anonymous):

\[x-1/x ^{2}-5\]

OpenStudy (ash2326):

Yeah, it's a cubic. Could you write what do you get?

OpenStudy (anonymous):

how to solve cubic equation!

OpenStudy (anonymous):

=\[x/(x ^{2}-5) -1/(x ^{2}-5)\]

OpenStudy (anonymous):

cut -1/x^2-5 from both side

OpenStudy (anonymous):

hey u got my question wrong @mathslover plzz lokk it once again

OpenStudy (anonymous):

x - 1/(x^2-25)

OpenStudy (anonymous):

so u have x/x^2-5=5 x=5x^2-25

OpenStudy (anonymous):

@dg123 @mathslover x - 1/(x^2-25)

mathslover (mathslover):

oh wait

OpenStudy (anonymous):

ooooh

OpenStudy (anonymous):

then there is only 1 solution tht is 5

OpenStudy (anonymous):

the answer should be 0

OpenStudy (anonymous):

@ash2326 plzz help confused

OpenStudy (anonymous):

the answer is 0, yeh bcoz the solution u get is 5, but it cant be putted on the denominator or else it will make the denominator =0

OpenStudy (anonymous):

which aint happening

OpenStudy (anonymous):

plzz @dg123 can u show the stepssssss

OpenStudy (anonymous):

yeh i am giving

OpenStudy (anonymous):

x-(1/x^2-5)=5-(1/x^2-5) add 1/x^2-5 on both sides, u get x=5

OpenStudy (anonymous):

so according to u it is 1 but it should be 0

OpenStudy (anonymous):

but look at the term 1/x^2-25 if u put x=5, the denominaot turns 0

OpenStudy (maheshmeghwal9):

@Yahoo! would u write the cubic equation which u gt?

OpenStudy (anonymous):

but dividing by zero is not possible, so there is no solution

mathslover (mathslover):

i think that i did the answer @Yahoo! will that be helping u ?

OpenStudy (ash2326):

@Yahoo! when we cancel the terms from both side, we assume that x is not 5, but on solving we get x=5 so there is no real solution for this

OpenStudy (anonymous):

in a nutshell, ( we get one solutio which is 5 but it makes the terms undefined, so there is no solution)